# Prometheus Mind: Retrofitting Memory to Frozen Language Models

### v1

@@ Abstract
On PrometheusExtract-132 (132 cases), the system achieves 94.4% retrieval on clean inputs (n=54, 95% CI: [84.9%, 98.1%]), degrading to 19.4% on informal inputs with ellipsis, filler words, or implicit subjects (n=36). The primary bottleneck is relation classification (47.3% accuracy), responsible for most extraction errors.
@@ 1 Introduction
We propose a third path: memory as a removable module for frozen language models. Our system, Prometheus Mind, adds memory capabilities to a frozen Qwen3-4B model through 11 small adapter networks (530MB total, 7% parameter overhead) and hidden state injection hooks. Critically, unlike RAG which injects memory as context tokens, we inject memory directly into the model’s hidden state representations at 86% depth—the model processes the memory through its remaining layers rather than reading it as text. And unlike Titans, our approach requires no architectural changes: the base model weights remain frozen, and removing the adapters restores the original model exactly.
A key insight from our development is that simply training an adapter to inject memory is insufficient—the frozen model learns to ignore injected signals, achieving only 25% accuracy. The breakthrough came from recognizing that the model itself must learn to accept injections. By training lightweight LoRA adapters on the model’s attention layers while keeping the injection mechanism fixed, we create a “reception pathway” that enables the model to utilize injected memory, improving accuracy to 87.5%.
@@
The result: memory for a frozen Qwen3-4B model via 11 modular adapters (530MB, 7% overhead), achieving 94.4% retrieval on clean inputs (n=54, 95% CI: [84.9%, 98.1%]) and 19.4% on adversarial inputs (n=36). We evaluate on PrometheusExtract-132, a benchmark of 132 cases spanning 10 difficulty categories with 95% confidence intervals for all metrics.
@@ 4.1 System Overview
    Injection: The retrieved answer is injected into the model’s hidden states at 86% depth using the lm_head.weight as the value encoder, enabling the model to generate responses grounded in memory.
@@
Figure 1: Prometheus Mind architecture. Input flows through a frozen Qwen3-4B (left). CDD extracts facts into structured memory (right). At query time, multi-hop reasoning retrieves answers and injects them at layer 30 using Identity V (lm_head.weight).
@@ 4.3.3 Adapter Inventory
4.4 Memory Injection via QKV-Adapter Interface
@@ 4.4.2 The QKV-Adapter Breakthrough
We inject a memory key-value pair at layer 30 (86% depth) during the first generated token only. The model attends to this injected K-V pair through its normal attention computation, naturally incorporating the memory into its processing.
@@
We discovered that no training is needed. The model’s own lm_head.weight already contains the answer:
@@ 6.4 Limitations
    Injection layer sensitivity: The 86% depth injection point was empirically tuned; theoretical understanding is lacking.
@@ C.4 Injection Depth Sensitivity
We evaluated injection at different layer depths:
Table 11: Accuracy by injection depth (layer / total layers).
Depth Layer Accuracy
50% 18/36 45%
70% 25/36 72%
86% 30/36 91%
92% 33/36 78%
97% 35/36 62%
The optimal injection point (86% depth) balances having enough processing capacity remaining while being late enough that the model has built sufficient context. Earlier injection gets diluted; later injection causes repetition.

### v2

@@ Abstract
On PrometheusExtract-132 (132 cases spanning 10 difficulty categories), the system achieves 94.4% retrieval on clean inputs (n=54, 95% CI: [84.9%, 98.1%]), 57.6% overall, and 19.4% on adversarial inputs with ellipsis, filler words, or implicit subjects (n=36). The primary bottleneck is relation classification (47.3% accuracy), responsible for most extraction errors.
@@ 1 Introduction
We propose a third path: memory as a removable module for frozen language models. Our system, Prometheus Mind, adds memory capabilities to a frozen Qwen3-4B model through 11 small adapter networks (530MB total, 7% parameter overhead) and hidden state injection hooks. Critically, unlike RAG which injects memory as context tokens, we inject memory directly into the model’s attention mechanism at layer 35 of 36 (97% depth)—the model attends to the memory through its native attention computation rather than reading it as text. And unlike Titans, our approach requires no architectural changes: the base model weights remain frozen, and removing the adapters restores the original model exactly.
A key insight from our development is that simply training an adapter to inject memory is insufficient—the frozen model learns to ignore injected signals. We initially explored LoRA adapters on attention layers to create a “reception pathway,” achieving 87.5% accuracy. However, our final breakthrough was simpler: by using the model’s own lm_head.weight rows as value vectors (Identity V), the frozen model generates correct outputs without modifying base model weights. Unlike learned value encoders that memorize training examples and fail on unseen facts, Identity V generalizes to any token in the vocabulary because it uses the model’s native output representations.
@@
The result: memory for a frozen Qwen3-4B model via 11 modular adapters (530MB, 7% overhead), achieving 94.4% retrieval on clean inputs (n=54, 95% CI: [84.9%, 98.1%]), 57.6% overall, and 19.4% on adversarial inputs (n=36). We evaluate on PrometheusExtract-132, a benchmark of 132 cases spanning 10 difficulty categories with 95% confidence intervals for all metrics.
@@ 4.1 System Overview
    Injection: The retrieved answer is injected into the model’s the attention mechanism at layer 35 (97% depth) using lm_head.weight as the value encoder, enabling the model to generate responses grounded in memory.
@@
Figure 1: Prometheus Mind architecture. Input flows through a frozen Qwen3-4B (left). CDD extracts facts into structured memory (right). At query time, multi-hop reasoning retrieves answers and injects them at layer 35 via attention K-V pairs, using Identity V (lm_head.weight).
@@ 4.3.3 Adapter Inventory
4.4 Memory Injection via Attention K-V Pairs
@@ 4.4.2 The Attention K-V Breakthrough
We inject a memory key-value pair at layer 35 (97% depth) during the first generated token only. The model attends to this injected K-V pair through its normal attention computation, naturally incorporating the memory into its processing.
Why first-token-only?
Without this constraint, the model enters repetition loops (“Mark Mark Mark…”). By injecting only on the first generated token, the model outputs the injected value once, then continues naturally. This was a critical discovery during development.
@@
We discovered that no training is needed for V. The model’s own lm_head.weight already contains the answer. (The K encoder—a small MLP mapping token embeddings to attention key space—does require training, but V is entirely training-free.)
@@ 6.4 Limitations
    Injection layer sensitivity: The layer 35 (97% depth) injection point was empirically tuned; theoretical understanding is lacking.
  6. 6.
    Single-token injection: Current implementation injects one token at a time. Multi-token answers (e.g., “New York”) require extension.
@@ 7 Conclusion
Broader Implications.
While demonstrated for memory, these principles may generalize to any capability expressible as routing to existing representations. The frozen model contains latent capabilities; adapters provide access patterns. This suggests a paradigm of modular cognition: composable, reversible capability modules over a fixed base. The value of reversibility extends beyond convenience—it enables A/B testing, rollback, and compliance with deployment constraints that prohibit weight modification.
@@ C.4 Injection Depth Sensitivity
We evaluated injection at different layer depths using two methods:
Hidden state blending (early experiments).
Direct addition of memory embeddings to hidden states showed optimal performance at 86% depth (layer 30/36), with degradation at both earlier and later layers. However, this method achieved only 91% accuracy even at optimal depth.
Attention K-V injection (final method).
The breakthrough came from injecting memory as key-value pairs into the attention mechanism at layer 35 (97% depth). Combined with first-token-only injection and Identity V, this achieves near-perfect retrieval accuracy. The key insight: at 97% depth, the model has built full context, and attention-based injection avoids the repetition issues that plagued direct hidden state modification at late layers.
Table 11: Accuracy by injection method and depth.
Method Depth Layer Accuracy
Hidden state blend 86% 30/36 91%
Hidden state blend 97% 35/36 62% (repetition)
Attention K-V 97% 35/36 $\sim$100%
The difference illustrates that how memory is injected matters as much as where. Attention-based injection at late layers succeeds where direct hidden state modification fails.